40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Sat Oct 26 21:46:33 2013, RANA, Update, IOO, Mode Cleaner Tune-UP MCtrend.pdfMC40days.png
    Reply  Mon Oct 28 14:12:48 2013, manasa, Update, IOO, Mode Cleaner relocked Screenshot-Untitled_Window-3.png
       Reply  Wed Oct 30 01:53:52 2013, Jenne, Update, IOO, Mode Cleaner relocked 
    Reply  Mon Oct 28 21:33:55 2013, RANA, Update, IOO, Mode Cleaner Tune-UP Untitled.png
       Reply  Thu Oct 31 15:34:28 2013, manasa, Update, IOO, MC not happy since yesterday MC_drift.png
       Reply  Thu Oct 31 20:05:48 2013, RANA, Update, IOO, Mode Cleaner Tune-UP drifty.png
          Reply  Fri Nov 1 09:45:32 2013, Steve, Update, IOO, beam dumps to be find razor_beam_dumps.jpg
          Reply  Mon Nov 4 11:37:12 2013, Steve, Update, IOO, PSL output shutter installed Shutterisback.jpg
             Reply  Mon Nov 4 12:59:43 2013, Jenne, Update, IOO, MC alignment not so good after PSL output shutter installed MCspots_4Nov2013.png
                Reply  Tue Nov 5 00:39:43 2013, manasa, Update, IOO, IFO alignment tuning MC_spot.png
                   Reply  Tue Nov 5 08:44:21 2013, Steve, Update, IOO, after last steering mirror mount swap after_Shutter_SMm_installed.pngafterlastSMmountswap.jpg
                   Reply  Tue Nov 5 17:12:34 2013, Steve, Update, LSC, centered qpds recentered.png
                      Reply  Wed Nov 6 08:33:53 2013, Steve, Update, IOO, Poiting changes of PSL 1dayTRENDofPOITING.png20dayTRENDioo.png
                         Reply  Thu Nov 7 08:57:20 2013, Steve, Update, IOO, PSL pointing monitoring PMC11072013.jpgPMCTqpdMon.png
                            Reply  Mon Nov 11 22:35:45 2013, RANA, Update, IOO, PSL pointing monitoring qpd.pdf
                               Reply  Tue Nov 12 23:48:23 2013, RANA, Update, IOO, PSL pointing monitoring 
                               Reply  Thu Nov 21 16:32:20 2013, Steve, Update, IOO, PMC needs a touch of love PMCplus.png
                               Reply  Fri Dec 6 15:57:41 2013, Steve, Update, IOO, beam dumps for PSL pointing monitoring  IOO_QPD_MONS.jpgPointingTrend23d.png
                                  Reply  Tue Dec 10 10:07:01 2013, Steve, Update, IOO, more beam traps InputWindowRefDump.jpgInpWindowRefDupm.jpg
Message ID: 9342     Entry time: Tue Nov 5 00:39:43 2013     In reply to: 9336     Reply to this: 9343   9347
Author: manasa 
Type: Update 
Category: IOO 
Subject: IFO alignment tuning 

Information acknowledged from Steve:

The last steering mirror mount for IR on the PSL was swapped for a more robust one. Prior to swapping the ibeam positions on the PSL IOO QPDS in ang and pos were recorded.

What I did henceforth:

1. Once the last steering mirror was installed, I walked the beam to restore input pointing using the last 2 steering mirrors. It needed a lot of work in yaw as expected.

2. When the input pointing was recovered, MC locked right away in TEM00. I measured the MC spot positions and compared it with Jenne's measurement made prior to the swap. The spot positions were pretty close.

3. The input pointing was adjusted in pitch and yaw (on the last steering mirror) in small steps. MC alignment was recovered and spot positions were measured each time. After several iterations, the MC spot positions were pretty much centered. I recentered the WFS and reset the WFS offsets. MC is now locked with WFS enabled at ~16900 counts.

MC_spot.png

4. Since the arms were aligned this morning, I used the Y arm as reference and corrected for the input pointing using tip-tilts.

5. Arms locked right away. Note: ASS doesn't seem to be doing it's job. I had to manually align the arms for maximum on TRX and TRY.

6. MICH and PRMI lock were also recovered.

7. I started to check the status with ALS as well. But for reasons unknown, I don't see any ADC counts corresponding to the beat note. Looking at the beatbox there aren't any signs of disconnected cables.  I am saving this as a morning job to fix it.

ELOG V3.1.3-